OpenStudy (anonymous):

consider a sequence given by the formula f(n)=13(n) -1 starting with n=1. Generate the first 5 terms of the sequence. Please help. Thank you so much

OpenStudy (anonymous):

Assuming that n increments by 1, all you need to do is substitute n for values 1, 2, 3, 4, 5.

OpenStudy (anonymous):

Example: f(1) = 13(1) - 1 f(1) = 13 - 1 f(1) = 12

OpenStudy (anonymous):

would the next be 25?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

thank you. I thought it was harder than this. so easy. you are so nice bye